The most recent financial statements for Moose Tours, Inc., appear below. Sales for 2016 are projected to grow by 20 percent. Interest expense will remain constant; the tax rate and the dividend payout rate will also remain constant. Costs, other expenses, current assets, fixed assets, and accounts payable increase spontaneously with sales.

MOOSE TOURS, INC.
2015 Income Statement
  Sales $ 749,000
  Costs 584,000
  Other expenses 20,000
  Earnings before interest and taxes $ 145,000
  Interest expense 16,000
  Taxable income $ 129,000
  Taxes (30%) 38,700
  Net income $ 90,300
   
    Dividends $ 18,060
    Addition to retained earnings 72,240

  

MOOSE TOURS, INC.
Balance Sheet as of December 31, 2015
Assets Liabilities and Owners’ Equity
  Current assets   Current liabilities
    Cash $ 20,840     Accounts payable $ 55,000
    Accounts receivable 33,160     Notes payable 14,200
    Inventory 70,120
      Total $ 69,200
      Total $ 124,120   Long-term debt $ 132,000
  Fixed assets   Owners’ equity
    Net plant and equipment $ 420,000     Common stock and paid-in surplus $ 118,000
    Retained earnings 224,920
      Total $ 342,920
  Total assets $ 544,120   Total liabilities and owners’ equity $ 544,120

If the firm is operating at full capacity and no new debt or equity is issued, what external financing is needed to support the 20 percent growth rate in sales? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
Answer #1

External Funds Needed = (A0/S0*(S1-S0)) - (L0/S0*(S1-S0)) - (PM*S1*b) - New common stock

where

Ao - Assets (at time 0) which vary directly with Sales = 544120

Lo - Liabilities (at time 0) which vary directly with Sales=55000

So - Current Sales = 749000

S1 - Projected Sales = 749000*1.32= 898800

b - Retention ratio = 72240/90300 = 80%

PM - Profit Margin = 90300/749000

External Funds Needed = (544120/749000*(898800-749000)) - (55000/749000*(898800-749000)) - (90300/749000*898800*.8) - 0

= 108824- 11000 - 86688

= 11,136